# SimPO: Simple Preference Optimization with a Reference-Free Reward

### v1

@@ Abstract
Direct Preference Optimization (DPO) is a widely used offline preference optimization algorithm that reparameterizes reward functions in reinforcement learning from human feedback (RLHF) to enhance simplicity and training stability. In this work, we propose SimPO, a simpler yet more effective approach. The effectiveness of SimPO is attributed to a key design: using the average log probability of a sequence as the implicit reward. This reward formulation better aligns with model generation and eliminates the need for a reference model, making it more compute and memory efficient. Additionally, we introduce a target reward margin to the Bradley-Terry objective to encourage a larger margin between the winning and losing responses, further enhancing the algorithm’s performance. We compare SimPO to DPO and its latest variants across various state-of-the-art training setups, including both base and instruction-tuned models like Mistral and Llama3. We evaluated on extensive instruction-following benchmarks, including AlpacaEval 2, MT-Bench, and the recent challenging Arena-Hard benchmark. Our results demonstrate that SimPO consistently and significantly outperforms existing approaches without substantially increasing response length. Specifically, SimPO outperforms DPO by up to 6.4 points on AlpacaEval 2 and by up to 7.5 points on Arena-Hard. Our top-performing model, built on Llama3-8B-Instruct, achieves a remarkable 44.7 length-controlled win rate on AlpacaEval 2—surpassing Claude 3 Opus on the leaderboard, and a 33.8 win rate on Arena-Hard—making it the strongest 8B open-source model. [^0] Code and models can be found at https://github.com/princeton-nlp/SimPO.
@@ 1 Introduction
  • Simplicity: SimPO does not require a reference model making it more lightweight and easier to implement compared to DPO, and other reference-dependent approaches.
  • Significant performance advantage: Despite its simplicity, SimPO significantly outperforms DPO and its latest variants (e.g., a recent reference-free objective ORPO [38]). The performance advantage is consistnt across various training setups and extensive instruction-following benchmarks, including AlpacaEval 2 [51, 24] and the challenging Arena-Hard [50] benchmark. It achieves up to a 6.4 point improvement on AlpacaEval 2 and a 7.5 point improvement on Arena-Hard.
@@
Extensive analysis shows that SimPO utilizes preference data more effectively, leading to a more accurate likelihhod ranking of winning and losing responses on a held-out validation set, which in turn translates to a better policy model. As shown in Table 1, we produce a top-performing model, built on top of Llama3-8B-instruct, that achieves a notable 44.7 length-controlled win rate on AlpacaEval 2, outperforming Claude 3 Opus on the leaderboard, and a 33.8 win rate on Arena-Hard, making it the strongest 8B open-source model to date.
@@ 2.2 A Simple Reference-Free Reward Aligned with Generation
Direct maximization of this metric during decoding is intractable, and various decoding strategies can be used to approximate it, such as greedy decoding [30], beam search [31, 49], nucleus sampling [35], and top-$k$ sampling [26, 36, 61]. Additionally, this metric is commonly used for ranking options in multiple-choice tasks within language models [11, 37, 58]. In DPO, for any triple $(x,y_{w},y_{l})$, satisfying the reward ranking $r(x,y_{w})>r(x,y_{l})$ does not necessarily mean that the likelihood ranking $p_{\theta}(y_{w}\mid x)>p_{\theta}(y_{l}\mid x)$ is met. In fact, only roughly $50\%$ of the triples from a held-out set satisfy this condition when trained with DPO (see Figure 4(b)).
@@ 3 Experimental Setup
For the Instruct setup, we use off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct, or mistralai/Mistral-7B-Instruct-v0.2) as the SFT models. These models have undergone extensive instruction-tuning processes, making them more powerful and robust than the SFT models in the Base setup. However, they are also more opaque because their RLHF procedure is not publicly disclosed. To mitigate the distribution shift between SFT models and the preference optimization process, we generate the preference dataset using the SFT models following [74]. Specifically, we use prompts from the UltraFeedback dataset and regenerate the chosen and rejected response pairs $(y_{w},y_{l})$ with the SFT models. For each prompt $x$, we generate 5 responses using the SFT model with a sampling temperature of 0.8. We then use PairRM [41] to score the 5 responses, selecting the highest-scoring one as $y_{w}$ and the lowest-scoring one as $y_{l}$. For simplicity, we generate data in a single pass instead of iteratively as in [74].
@@
We compare SimPO with other offline preference optimization methods listed in Table 3.444Many recent studies [82, 70] have extensively compared DPO and PPO [66]. We will leave the comparison of PPO and SimPO to future work. IPO [6] is a theoretically grounded approach method that avoids DPO’s assumption that pairwise preferences can be replaced with pointwise rewards. KTO [25] learns from non-paired preference data. ORPO [38] introduces a reference-model-free odd ratio term to directly contrast winning and losing responses with the policy model and jointly trains with the SFT objective. R-DPO [60] is a modified version of DPO that includes an additional regularization term to prevent exploitation of length. We thoroughly tune the hyperparameters for each baseline and report best performance. We find that many variants of DPO do not empirically present an advantage over standard DPO. Further details can be found in Appendix A.
@@ 4.1 Main Results and Ablations
As shown in Table 4, while all preference optimization algorithms enhance performance over the SFT model, SimPO, despite its simplicity, achieves the best overall performance across all benchmarks and settings. Notably, SimPO outperforms the best baseline by 3.6 to 4.8 points on the AlpacaEval 2 LC win rate and by 0.2 to 6.2 points on Arena-Hard across various settings. These consistent and significant improvements highlight the robustness and effectiveness of SimPO.
@@
Although all three benchmarks are widely adopted, we find that MT-Bench exhibits poor separability across different methods. Minor differences between methods on MT-Bench may be attributed to randomness, likely due to the limited scale of its evaluation data and its single-instance scoring protocol. This finding aligns with observations reported in [50]. In contrast, AlpacaEval 2 and Arena-Hard provide more meaningful distinctions between different methods. Additionally, Arena-Hard employs a different judge model from the baseline model, likely leading to fairer evaluations. We observe that the win rate on Arena-Hard is significantly lower than on AlpacaEval 2, indicating that Arena-Hard is a more challenging benchmark.
@@ 4.3 The Impact of Target Reward Margin in SimPO
(a) Len. correlation.
(b) Contingency table.
@@
(d) Efficiency.
Figure 4: Comparison between SimPO and DPO, measured on UltraFeedback. (a) Spearman correlation between average log probability and response length for DPO. (b) Contingency table of rankings based on DPO rewards and the average log likelihood (measured on the training set). (c) Reward accuracy of DPO and SimPO. (d) Runtime and memory usage for DPO and SimPO.
@@
DPO reward implicitly facilitates length normalization.
Although the DPO reward expression $r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ (with the partition function excluded), lacks an explicit term for length normalization, the logarithmic ratio between the policy model and the reference model can serve to implicitly counteract length bias. As shown in Table 6 and Figure 4(a), employing DPO reduces the Spearman correlation coefficient between average log likelihood and response length compared to the approach without any length normalization (referred to as “SimPO w/o LN”). However, it still exhibits a stronger positive correlation when compared to SimPO. [^5] Note that this correlation does not fully reflect the generation length. Despite DPO showing a stronger correlation, the length of its generated responses is comparable to or even slightly shorter than those of the SimPO models. Please find more details in Appendix D.
DPO reward mismatches generation likelihood.
@@
Another benefit of SimPO is its efficiency as it does not use a reference model. Figure 4(d) illustrates the overall run time and per-GPU peak memory usage of SimPO and DPO in the Llama3-Base setting using 8×H100 GPUs. Compared to a vanilla DPO implementation, [^6] DPO can be as memory efficient as SimPO if it were implemented to separate the forward passes of the reference model from the actual preference optimization. However, this implementation is not standard practice. SimPO cuts run time by roughly 20% and reduces GPU memory usage by about 10%, thanks to eliminating forward passes with the reference model.
@@ 5 Related Work
Reinforcemant learning from human feedback.
@@
Preference optimization.
Given that online preference optimization algorithms are complex and difficult to optimize [89, 65], researchers have been exploring more efficient and simpler alternative offline algorithms. Direct Preference Optimization (DPO) [62] is a notable example. However, the absence of a reward model in DPO limits its ability to sample preference pairs from the optimal policy. To address this, researchers have explored augmenting preference data using a trained SFT policy [87] or a refined SFT policy with rejection sampling [55], enabling the policy to learn from data generated by the optimal policy. Further studies have extended this approach to an iterative training setup, continuously updating the reference model for optimization with either the most recent policy model or by generating preference pairs scored and selected by the evolving policy model [42, 63, 84]. Additionally, some studies have replaced the online RL objective with offline ranking objectives, allowing for comparisons among more than two instances from a broader perspective [23, 54, 68, 83]. Another line of work explores simpler preference optimization objectives without relying on a reference model [38, 81], similar to SimPO. Our work follows the pair-wise preference optimization setup without iterative training. We compare it to a series of offline algorithms, including DPO, IPO [6], KTO [25], ORPO [38], and R-DPO [60], and find that SimPO can outperform these methods in both efficiency and performance. Recently, [71] proposed a generalized preference optimization framework that unifies different offline algorithms, and SimPO can be seen as a special case of this framework.
@@ Acknowledgments
The authors would like to thank Tianyu Gao, Tanya Goyal, Kaifeng Lyu, Sadhika Malladi, Eric Mitchell, Lewis Tunstall and members of the Princeton NLP group for their valuable feedback and discussions. We thank Niklas Muennighoff for his advice on training and reproducing training KTO models. Mengzhou Xia is supported by an Apple Scholars in AIML Fellowship. This research is also funded by the National Science Foundation (IIS-2211779) and a Sloan Research Fellowship.
@@ Appendix A Implementation Details
KTO [25] $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1$

### v2

@@ Abstract
Direct Preference Optimization (DPO) is a widely used offline preference optimization algorithm that reparameterizes reward functions in reinforcement learning from human feedback (RLHF) to enhance simplicity and training stability. In this work, we propose SimPO, a simpler yet more effective approach. The effectiveness of SimPO is attributed to a key design: using the average log probability of a sequence as the implicit reward. This reward formulation better aligns with model generation and eliminates the need for a reference model, making it more compute and memory efficient. Additionally, we introduce a target reward margin to the Bradley-Terry objective to encourage a larger margin between the winning and losing responses, further enhancing the algorithm’s performance. We compare SimPO to DPO and its latest variants across various state-of-the-art training setups, including both base and instruction-tuned models like Mistral and Llama3. We evaluate on extensive instruction-following benchmarks, including AlpacaEval 2, MT-Bench, and the recent challenging Arena-Hard benchmark. Our results demonstrate that SimPO consistently and significantly outperforms existing approaches without substantially increasing response length. Specifically, SimPO outperforms DPO by up to 6.4 points on AlpacaEval 2 and by up to 7.5 points on Arena-Hard. Our top-performing model, built on Llama3-8B-Instruct, achieves a remarkable 53.7 length-controlled win rate on AlpacaEval 2—surpassing Claude 3 Opus on the leaderboard, and a 36.5 win rate on Arena-Hard—making it the strongest 8B open-source model. [^0] Code and models can be found at https://github.com/princeton-nlp/SimPO.
@@ 1 Introduction
  • Simplicity: SimPO does not require a reference model, making it more lightweight and easier to implement compared to DPO and other reference-dependent approaches.
  • Significant performance advantage: Despite its simplicity, SimPO significantly outperforms DPO and its latest variants (e.g., a recent reference-free objective ORPO [40]). The performance advantage is consistent across various training setups and extensive instruction-following benchmarks, including AlpacaEval 2 [53, 26] and the challenging Arena-Hard [52] benchmark. It achieves up to a 6.4 point improvement on AlpacaEval 2 and a 7.5 point improvement on Arena-Hard.
@@
Extensive analysis shows that SimPO utilizes preference data more effectively, leading to a more accurate likelihood ranking of winning and losing responses on a held-out validation set, which in turn translates to a better policy model. As shown in Table 1, we produce a top-performing model, built on top of Llama3-8B-instruct, that achieves a notable 44.7 length-controlled win rate on AlpacaEval 2, outperforming Claude 3 Opus on the leaderboard, and a 33.6 win rate on Arena-Hard, making it the strongest 8B open-source model to date. By training on a preference optimization dataset annotated by a stronger reward model (i.e., ArmoRM [80]), we further produce a stronger v0.2 model with an AlpacaEval 2 LC win rate of 53.7 and an Arena-Hard win rate of 36.5 (Appendix G).
@@ 2.2 A Simple Reference-Free Reward Aligned with Generation
Direct maximization of this metric during decoding is intractable, and various decoding strategies can be used to approximate it, such as greedy decoding [32], beam search [33, 51], nucleus sampling [37], and top-$k$ sampling [28, 38, 63]. Additionally, this metric is commonly used for ranking options in multiple-choice tasks within language models [12, 39, 60]. In DPO, for any triple $(x,y_{w},y_{l})$, satisfying the reward ranking $r(x,y_{w})>r(x,y_{l})$ does not necessarily mean that the likelihood ranking $p_{\theta}(y_{w}\mid x)>p_{\theta}(y_{l}\mid x)$ is met. In fact, only roughly $50\%$ of the triples from the training set satisfy this condition when trained with DPO (see Figure 4(b)).
@@ 3 Experimental Setup
For the Instruct setup, we use off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct, or mistralai/Mistral-7B-Instruct-v0.2) as the SFT models. [^2] It is unclear whether the released instruct checkpoints have undergone supervised fine-tuning (SFT) or the complete reinforcement learning from human feedback (RLHF) pipeline. For simplicity, we refer to these checkpoints as SFT models. These models have undergone extensive instruction-tuning processes, making them more powerful and robust than the SFT models in the Base setup. However, they are also more opaque because their RLHF procedure is not publicly disclosed. To mitigate the distribution shift between SFT models and the preference optimization process, we generate the preference dataset using the SFT models following [76]. This makes our Instruct setup closer to an on-policy setting. Specifically, we use prompts from the UltraFeedback dataset and regenerate the chosen and rejected response pairs $(y_{w},y_{l})$ with the SFT models. For each prompt $x$, we generate 5 responses using the SFT model with a sampling temperature of 0.8. We then use llm-blender/PairRM [43] to score the 5 responses, selecting the highest-scoring one as $y_{w}$ and the lowest-scoring one as $y_{l}$. For simplicity, we generate data in a single pass instead of iteratively as in [76]. In addition, we create a v0.2 Llama3-Instruct setting by using RLHFlow/ArmoRM-Llama3-8B-v0.1 [80] as the reward model to rank generated data, which yields significantly improved performance (more details in Appendix G).
@@
RRHF [87] $\max\left(0,-\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)+\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$
SLiC-HF [92] $\max\left(0,\delta-\log\pi_{\theta}(y_{w}|x)+\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$
@@
CPO [84] $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$
@@
We compare SimPO with other offline preference optimization methods listed in Table 3.555Many recent studies [86, 72] have extensively compared DPO and PPO [68]. We will leave the comparison of PPO and SimPO to future work. RRHF [87] and SLiC-HF [92] are ranking losses. RRHF uses length-normalized log-likelihood, similar to SimPO’s reward function, while SLiC-HF uses log-likelihood directly and includes an SFT objective. IPO [6] is a theoretically grounded approach method that avoids DPO’s assumption that pairwise preferences can be replaced with pointwise rewards. CPO [84] uses sequence likelihood as a reward and trains alongside an SFT objective. KTO [27] learns from non-paired preference data. ORPO [40] [^6] ORPO can directly train on preference data without the SFT stage. For fair comparisons, we start ORPO from the same SFT checkpoints as other baselines, which should yield better results than starting from a base checkpoint. introduces a reference-model-free odd ratio term to directly contrast winning and losing responses with the policy model and jointly trains with the SFT objective. R-DPO [62] is a modified version of DPO that includes an additional regularization term to prevent exploitation of length. We thoroughly tune the hyperparameters for each baseline and report best performance. We find that many variants of DPO do not empirically present an advantage over standard DPO. Further details can be found in Appendix A.
@@ 4.1 Main Results and Ablations
As shown in Table 4, while all preference optimization algorithms enhance performance over the SFT model, SimPO, despite its simplicity, achieves the best overall performance across all benchmarks and settings. These consistent and significant improvements highlight the robustness and effectiveness of SimPO. Notably, SimPO outperforms the best baseline by 3.6 to 4.8 points on the AlpacaEval 2 LC win rate across various settings. On Arena-Hard, SimPO consistently achieves superior performance, though it is occasionally surpassed by CPO [84]. We find that CPO generates responses that are, on average, 50% longer than those generated by SimPO (See Table 10). Arena-Hard might favor longer generations due to the absence of a length penalty in its evaluation.
@@
Although all three benchmarks are widely adopted, we find that MT-Bench exhibits poor separability across different methods. Minor differences between methods on MT-Bench may be attributed to randomness, likely due to the limited scale of its evaluation data and its single-instance scoring protocol. This finding aligns with observations reported in [52]. In contrast, AlpacaEval 2 and Arena-Hard provide more meaningful distinctions between different methods. We observe that the win rate on Arena-Hard is significantly lower than on AlpacaEval 2, indicating that Arena-Hard is a more challenging benchmark. [^7] Although our models excel on benchmarks, these evaluations have limitations, including restricted query space and potential biases from model-based evaluations. Efforts like WildBench [89] aim to expand these spaces, where SimPO models demonstrate competitive performance.
@@ 4.3 The Impact of Target Reward Margin in SimPO
(a) Length correlation (DPO).
(b) Contingency table (DPO).
@@
Figure 4: Comparison between SimPO and DPO, measured on UltraFeedback. (a) Spearman correlation between average log probability and response length for DPO. (b) Contingency table of rankings based on DPO rewards and the average log likelihood (measured on the training set). (c) Reward accuracy of DPO and SimPO.
(a) KL divergence w/ different $\beta$.
(b) Performance w/ different $\beta$.
(c) Efficiency of DPO vs. SimPO.
Figure 5: Comparison between SimPO and DPO (continued). (a) With different $\beta$ in DPO and SimPO, KL divergence from the policy model to the reference model on winning responses. (b) AlpacaEval2 LC win rate of DPO and SimPO with different $\beta$. (c) Runtime and memory usage for DPO and SimPO.
@@
DPO reward implicitly facilitates length normalization.
Although the DPO reward expression $r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ (with the partition function excluded) lacks an explicit term for length normalization, the logarithmic ratio between the policy model and the reference model can serve to implicitly counteract length bias. As shown in Table 6 and Figure 4(a), employing DPO reduces the Spearman correlation coefficient between average log likelihood and response length compared to the approach without any length normalization (referred to as “SimPO w/o LN”). However, it still exhibits a stronger positive correlation when compared to SimPO. [^8] Note that this correlation does not fully reflect the generation length. Despite DPO showing a stronger correlation, the length of its generated responses is comparable to or even slightly shorter than those of the SimPO models. Please find more details in Appendix D.
DPO reward mismatches generation likelihood.
@@
KL divergence of SimPO and DPO.
In Figure 5(a), we present the KL divergence between the policy model trained with DPO and SimPO and the reference model with different $\beta$, measured on the winning responses from a held-out set during training. Figure 5(b) shows the corresponding AlpacaEval 2 LC win rate. Although SimPO does not apply any form of regularization against the reference model, the KL divergence of SimPO is reasonably small. Increasing $\beta$ reduces the KL divergence for both DPO and SimPO, with DPO exhibiting a more pronounced reduction at higher $\beta$ values. In this particular setting (Mistral-base), Figure 5(b) demonstrates that a smaller $\beta$ can yield better AlpacaEval 2 performance, despite the higher KL divergence. [^9] We observe that in some settings (e.g., Llama3-Instruct), having a large $\beta$ (e.g., $\beta=10$) leads to better performance. We hypothesize that when the reference model is weak, strictly constraining the policy model to the reference model may not be beneficial. As a caveat, while we did not observe any training collapse or degeneration with proper tuning, in principle, training with SimPO could potentially lead to reward hacking since SimPO does not regularize against the reference model. In such a scenario, the model might achieve a low loss but degenerate in sequence generation.
@@
Another benefit of SimPO is its efficiency as it does not use a reference model. Figure 5(c) illustrates the overall run time and per-GPU peak memory usage of SimPO and DPO in the Llama3-Base setting using 8×H100 GPUs. Compared to a vanilla DPO implementation, [^10] DPO can be as memory efficient as SimPO if it were implemented to separate the forward passes of the reference model from the actual preference optimization. However, this implementation is not standard practice. SimPO cuts run time by roughly 20% and reduces GPU memory usage by about 10%, thanks to eliminating forward passes with the reference model.
@@ 5 Related Work
Reinforcement learning from human feedback.
@@
Offline vs. iterative preference optimization.
Given that online preference optimization algorithms are complex and difficult to optimize [95, 67], researchers have been exploring more efficient and simpler alternative offline algorithms. Direct Preference Optimization (DPO) [64] is a notable example. However, the absence of an explicit reward model in DPO limits its ability to sample preference pairs from the optimal policy. To address this, researchers have explored augmenting preference data using a trained SFT policy [92] or a refined SFT policy with rejection sampling [57], enabling the policy to learn from data generated by the optimal policy. Further studies have extended this approach to an iterative training setup, continuously updating the reference model for optimization with either the most recent policy model or by generating preference pairs scored and selected by the evolving policy model [44, 65, 88, 25, 83]. In this work, we focus exclusively on offline settings, avoiding any iterative training processes.
Preference optimization objectives.
A variety of preference optimization objectives have been proposed besides DPO. Ranking objectives allow for comparisons among more than two instances [24, 56, 70, 87]. Another line of work explores simpler preference optimization objectives that do not rely on a reference model [40, 85], similar to SimPO. [8] proposes a method to jointly optimize instructions and responses, finding it effectively improves DPO. [93] focuses on post-training extrapolation between the SFT and the aligned model to further enhance model performance. In this work, we compare SimPO to a series of offline algorithms, including RRHF [87], SLiC-HF [92], DPO [64], IPO [6], CPO [84], KTO [27], ORPO [40], and R-DPO [62], and find that SimPO can outperform these methods in both efficiency and performance. Recently, [73] proposed a generalized preference optimization framework that unifies different offline algorithms, and SimPO can be seen as a special case of this framework.
@@ Acknowledgments
The authors would like to thank Li Dong, Tianyu Gao, Tanya Goyal, Di Jin, Yuchen Lin, Kaifeng Lyu, Sadhika Malladi, Eric Mitchell, Lewis Tunstall, Haoxiang Wang, Wei Xiong, Zhen Xu, Libing Yang, Zhiyu Zhao, and members of the Princeton NLP group for their valuable feedback and discussions. We thank Niklas Muennighoff for his advice on training and reproducing training KTO models. We thank Haoran Xu for helping verify our CPO runs. Mengzhou Xia is supported by an Apple Scholars in AIML Fellowship. This research is also funded by the National Science Foundation (IIS-2211779) and a Sloan Research Fellowship.
@@ Appendix A Implementation Details
RRHF [87] $\max\left(0,-\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)+\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\lambda\in[0.1,0.5,1.0,10.0]$
SLiC-HF [92] $\max\left(0,\delta-\log\pi_{\theta}(y_{w}|x)+\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\lambda\in[0.1,0.5,1.0,10.0]$
$\beta\in[0.1,0.5,1.0,2.0]$
@@
CPO [84] $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
KTO [27] $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1.0$